OpenStudy (anonymous):

Evaluate the expression. r = <5, 9, -4>, v = <3, 3, -8>, w = <9, -6, -1> v ⋅ w

OpenStudy (freckles):

let x=<a,b,c> and y=<d,e,f> then x dot y=ad+be+fc

OpenStudy (freckles):

multiply first components multiply second components multiply third components then all the results of those products

OpenStudy (anonymous):

<27,-18,8>

OpenStudy (anonymous):

is that correct ? @freckles

OpenStudy (freckles):

dot product isn't a vector

OpenStudy (freckles):

it is just a scalar

OpenStudy (freckles):

you need to add the products together

OpenStudy (anonymous):

17

OpenStudy (anonymous):

??

OpenStudy (freckles):

yeah \[v \cdot w=27-18+8=9+8=17\]

OpenStudy (anonymous):

thanks again :)

OpenStudy (freckles):

np